3. A particle of mass 6 kg is on a slope at an angel of 20o
to the horizontal. The coefficient of
friction between the particle and the slope is 0.1. The particle is 5 m from the bottom of the
slope. It is projected up the slope with a speed 4 m s-1
?
(a) Find the distance travelled up the slope from the starting point until the particle comes
to rest.
(b) Find the time until the particle reaches the bottom of the slope.
1
Expert's answer
2021-09-08T16:14:23-0400

"\text{Let's denote the X-axis along the slope}"

"\text{}""\text{Take the positive direction up to the top }"

"\text{of the inclined plane}"

"\text{We choose the Y- axis perpendicular to the X-axis}"

"\alpha=20\degree"

"m= 6kg"

"g= 9.8\frac{m}{s^2}"

"\mu = 0.1"

"v_0= 4\frac{m}{s}"

"h_0= 5m"

"(a)"

"\text{Projection of forces on an Y-axis:}"

"N-mg\cos \alpha= 0"

"N=mg\cos \alpha= 6*9.8*\cos20 \degree=55.25N"

"\text{Projection of forces on an X-axis:}"

"F_{fr}= \mu N= 55.25*0.1=5.53N"

"F= - F{fr}-mg\sin\alpha= -5.53-6*9.8*\sin20 \degree=-25.65N"

"F = ma"

"a=\frac{F}{m}=\frac{-25.65}{6}=-4.28\frac{m}{s^2}"

"v = v_0+at"

"v_1 =0\text{ the particle has stopped moving}"

"t = -\frac{v_0}{a}=\frac{4}{4.28}=0.93s"

"S = v_0t+\frac{at^2}{2}= 4*0.93-\frac{4.28*0.93^2}{2}=1.87m(1)"

"\text{Answer: }1.87m"

"(b)"

"\text{Projection of forces on an Y-axis:}"

"N-mg\cos \alpha= 0"

"N=mg\cos \alpha= 6*9.8*\cos20 \degree=55.25N"

"\text{Projection of forces on an X-axis:}"

"F_{fr}= \mu N= 55.25*0.1=5.53N"

"F= -F{fr}+mg\sin\alpha=- 5.53+6*9.8*\sin20 \degree=14.58N"

"F = ma"

"a=\frac{F}{m}=\frac{14.58}{6}=2.43\frac{m}{s^2}"

"v_0=0"

"S _1= \frac{at^2}{2}"

"t = \sqrt{\frac{2S_1}{a}}"

"S _1= S(\text{from (1))}+\frac{h_0}{\sin\alpha}=1.87+14.2=16.07m"

"t = \sqrt{\frac{2S_1}{a}}= \sqrt{\frac{2*16.07}{2.43}}=3.64s"

"\text{Answer: }3.64s"